OpenStudy (anonymous):

find the volume of a regular pyramid with an equilateral triangular base with side 16.8 inches and height 25.5 inches.

OpenStudy (anonymous):

would this be right? A=1/2(25.5)(67.2) =1/2(1713.6) = 856.8 in squared

OpenStudy (anonymous):

formula i found is volume = (Area * Height)/3 so... ((16.8*16.8)/2 * (25.5))/3

OpenStudy (anonymous):

so, i get... 1199.52 in cubed (volume is cubed, area squared)

OpenStudy (anonymous):

ok so i used the wrong formula then...thanks for pointing that out :)

OpenStudy (anonymous):

so i need to find the base area first right? before i can find the surface area.

OpenStudy (anonymous):

before you can find the volume, yeah

OpenStudy (anonymous):

\[the volum = base area Xthe heigt ; base area = 0.5 \times \sin (6o) \times (base side)^{2}\]